Assume that the equation xyz = 1 holds in a group G. Does

it follow that yzx = 1? That yxz = 1? Justify your answer.


1
Expert's answer
2022-06-21T12:24:10-0400

"xyz=1" implies that "x(yz)=1". Let "yz=a". Then we have "xa=1" and so "ax=1" since "a" is invertible and "a^{-1}=x". It follows that "(yz)x=1". Hence "yzx=1".


On the other hand, if "xyz=1" it is not always true that "yxz=1". To see that, let "G" be the group of "2\times2" real matrices and let "x=\left(\begin{smallmatrix}1&2\\0&2\end{smallmatrix}\right)", "y=\left(\begin{smallmatrix}0&1\\2&1\end{smallmatrix}\right)" and "z=\left(\begin{smallmatrix}-\frac{1}{2}&\frac34\\1&-1\end{smallmatrix}\right)". Then "xyz=\left(\begin{smallmatrix}1&0\\0&1\end{smallmatrix}\right)=1" in "G". But "yxz=\left(\begin{smallmatrix}2&-2\\5&-\frac92\end{smallmatrix}\right)\ne1".